# Thompson Sampling for Repeated Newsvendor

### v1

@@ 5. Extensions to Online Learning with Censored Feedback
In this section, we extend the regret analysis of TS for the repeated newsvendor problem to a broader class of online learning algorithms. We consider a setting where the demand $D_{t}$ in period $t$ is drawn from Weibull distribution parameterized by $\theta_{\star}$. The decision-maker selects an action $A_{t}$ in each period, resulting in an observed feedback of $\min\{D_{t},A_{t}\}$. The loss incurred in each period is defined as $l(A_{t})=\min\{D_{t},A_{t}\}$. The cumulative regret over $T$ periods is defined as:
$\operatorname{Regret}(T,\theta_{\star})=\sum_{t=1}^{T}\left(l(A_{t})-l(A_{\star})\right),$
where $A_{\star}$ denotes the optimal action that minimizes the expected loss, given by:
$A_{\star}=\arg\min_{A}\mathbb{E}_{D\sim\theta_{\star}}[l(A)].$
5.1. Key Assumptions and Results
@@
(10) $\left|\mathbb{E}[A_{t}]-A_{\star}\right|\leq C_{2}\left|\mathbb{E}[\frac{1}{\theta_{t}}]-\frac{1}{\theta_{\star}}\right|,$
@@
Assumption 2 is satisfied in the repeated newsvendor model. Specifically, Lemma 3.2 shows that the optimal action in the newsvendor problem is given by $y_{t}(\theta_{t})=\frac{1}{\theta_{t}}\left(-\ln(\frac{h}{p+h}\right)^{1/k}$. Let us show how the the conclusions in key step 2 and key step 3 hold under this assumption.
For the key step 2, the Lipschitz continuity assumption in (10) directly leads to (8). Additionally, Lemma 3.4 provides a generic estimation result for censored feedback under the Weibull distribution that is independent of the loss function or algorithm in use. Consequently, the conclusion of key step 2 (Section 3.2.2) holds under Assumption 2.
For key step 3, we observe that the lower bounds for $\frac{1}{\theta_{t}}$ (as shown in inequalities (9b) to (9d)) are general results for censored feedback under the Weibull distribution and do not depend on the loss function or the specific algorithm used. Therefore, as long as the positive function $y_{t}$ is a monotone in $\frac{1}{\theta_{t}}$, a uniform lower bound for $y_{t}$ is guaranteed.
By synthesizing the above analysis on how the conclusions of all three key steps hold, we establish the following theorem on cumulative regret:
Theorem 5.1 (Regret of TS for general online learining with censored feedback).
@@
5.2. Technical limitation, possible refinement, and open questions
We highlight a technical limitation in Assumptions 1 and 2, which we believe can be addressed with additional complexity, as well as a more challenging open question.
First, the Lipschitz continuity assumptions currently involve expectations inside the absolute value. A more natural formulation would be to remove the expectation from these assumptions (e.g., by moving it outside the absolute value). This adjustment can be justified if the distribution of $A_{t}$ exhibits concentration properties (e.g., light-tailed, sub-Gaussian, etc.), though it would introduce additional complexity through high-probability arguments and tail assumptions.
Second, a more significant challenge lies in relaxing the Weibull distribution assumption. If successful, this would represent a substantial step forward from the current analysis. In particular, it would allow Assumption 2 to be stated directly in terms of $\theta$ rather than $\frac{1}{\theta}$, making it conceptually more natural. We conclude that relaxing the Weibull assumption remains a central open question in developing a more general theory for online learning with censored feedback, especially in higher-dimensional action spaces. We leave this extension for future work.
6. Conclusions
We present the first systematic study on applying Thompson Sampling (TS) to the repeated newsvendor problem and provide an initial exploration of how our analytical framework can be extended to broader online learning problems with censored feedback. We establish frequentist regret bounds and offer insights into how TS automatically balances the trade-off between “large exploration” and “optimal exploitation.” Our analysis follows three key steps, which naturally generalize to broader settings.
This work opens up a range of compelling research directions. A key avenue for future exploration is extending regret analysis to broader online learning environments with censored feedback, particularly by relaxing the Weibull demand assumption to develop a more flexible and general framework. Additionally, applying TS to broader economic settings—such as auctions, dynamic pricing, and real-time resource allocation—presents exciting opportunities, as willingness-to-pay observations are often censored in these contexts. Advancing research in these areas has the potential to enhance decision-making under uncertainty, fostering more robust, efficient, and adaptive learning mechanisms for complex real-world problems.
Appendix A Appendix
A.1. Proof for Lemma 3.3
@@ Appendix A Appendix
The proof largely follows Lemma B2, B3 in (Chuang and Kim, 2023). We denote $H=\left\{H_{t}\right\}$ the natural filtration generated by the right-censored sales data, i.e $H_{t}=\sigma\left\{(Y_{i},\delta_{i}):i\leq t\right\}$, where $Y_{t}=D_{t}\wedge y_{t}$ and $\delta_{t}=1\left[D_{t}<y_{t}\right]$.
According to the proof of Lemma B2 and B3 in (Chuang and Kim, 2023), we have
@@
From (Chuang and Kim, 2023), we have
@@
Since $\theta_{t}\sim\operatorname{Gamma}(\alpha_{t},\beta_{t})$, we have $\frac{1}{\theta_{t}}\sim\operatorname{InverseGamma}(\alpha_{t},\beta_{t})$. According to (Chen, 2014) Theorem 20, we have
@@
Recall that $M_{t}=\sum_{i=0}^{t-1}\left(\delta_{i}-\mathbb{E}\left[\delta_{i}\mid\mathcal{H}_{i-1}\right]\right)$ defined in Lemma 3.5 and $M_{t}$ is a martingale with bounded increments (specifically, bounded by 2 ), by Azuma’s inequality we have,
@@
A.6. Auxiliary Lemmas
Lemma A.1.
@@ A.6. Auxiliary Lemmas
(11a) $\alpha_{t}-1$ $=\alpha_{0}+\sum_{i=0}^{t-1}\left(1-e^{-\theta_{\star}y_{i}^{k}}\right)+M_{t}-1$
(11b) $\geq\alpha_{0}-1+\left(t-1\right)\left(1-\exp\{-\theta_{\star}L^{k}\right)+M_{t}$
(11c) $\geq t\left(1-\exp\{-\theta_{\star}L^{k}\}\right)+\alpha_{0}-1-\left(1-\exp\{-\theta_{\star}L\}^{k}\right)-\sqrt{8t}\ln{\left(\frac{2t^{2}}{\delta}\right)}$
@@
(11a) is derived from Lemma 3.5. (11b) comes from the fact that $y_{t}\geq L$ for all $t$ when the event $\xi$ holds. (11c) comes from the following Lemma A.2, which is proved in Appendix A.5.
Lemma A.2.
@@
(12a) $\geq 1-\sum_{t=1}^{T}\frac{\delta}{T}-\sum_{t=1}^{T}\frac{\delta}{t^{2}}-\sum_{t=1}^{T}\frac{\delta}{t^{2}}-\sum_{t=1}^{T}\left(\frac{2}{e}\right)^{\alpha_{t}}$
(12b) $\geq 1-\delta-\frac{\pi^{2}}{6}\delta-\frac{\pi^{2}}{6}\delta-\delta$
@@
For (12a), the first term comes from Lemma 3.3, the second term comes from Lemma 3.4. The third term comes from Lemma A.2. The fourth term comes from Lemma 3.6. (12b) comes from the following, recall $\alpha_{0}\geq\frac{\ln{\frac{T}{\delta}}}{\ln{\frac{e}{2}}}$.

### v2

@@ A.1 Extensions to Online Learning with Censored Feedback
In this section, we extend the regret analysis of TS for the repeated newsvendor problem to a broader class of online learning algorithms. We consider a setting where the demand $D_{t}$ in period $t$ is drawn from Weibull distribution parameterized by $\theta_{\star}$. The decision-maker selects an action $A_{t}$ in each period, resulting in an observed feedback of $\min\{D_{t},A_{t}\}$. The loss incurred in each period is defined as $l(A_{t})=\min\{D_{t},A_{t}\}$. The cumulative regret over $T$ periods is defined as $\operatorname{Regret}(T,\theta_{\star})=\sum_{t=1}^{T}\left(l(A_{t})-l(A_{\star})\right)$, where $A_{\star}$ denotes the optimal action that minimizes the expected loss, given by $A_{\star}=\arg\min_{A}\mathbb{E}_{D\sim\theta_{\star}}[l(A)]$.
A.1.1 Key Assumptions and Results
@@
$\left|\mathbb{E}[A_{t}]-A_{\star}\right|\leq C_{2}\left|\mathbb{E}[\frac{1}{\theta_{t}}]-\frac{1}{\theta_{\star}}\right|,$ (7)
@@
Assumption 5 is satisfied in the repeated newsvendor model. Specifically, Lemma 3 shows that the optimal action in the newsvendor problem is given by $y_{t}(\theta_{t})=\frac{1}{\theta_{t}}\left(-\ln(\frac{h}{p+h}\right)^{1/k}$. Let us show how the the conclusions in key step 2 and key step 3 hold under this assumption.For the key step 2, the Lipschitz continuity assumption in (7) directly leads to (5). Additionally, Lemma 5 provides a generic estimation result for censored feedback under the Weibull distribution that is independent of the loss function or algorithm in use. Consequently, the conclusion of key step 2 (Section 3.2) holds under Assumption 5. For key step 3, we observe that the lower bounds for $\frac{1}{\theta_{t}}$ (as shown in inequalities (LABEL:lb-yt-b) to (2c)) are general results for censored feedback under the Weibull distribution and do not depend on the loss function or the specific algorithm used. Therefore, as long as the positive function $y_{t}$ is a monotone in $\frac{1}{\theta_{t}}$, a uniform lower bound for $y_{t}$ is guaranteed. By synthesizing the above analysis on how the conclusions of all three key steps hold, we establish the following theorem on cumulative regret:
Theorem 3 (Regret of TS for general online learining with censored feedback).
@@
A.2 Proof for Lemma 4
@@ A.3 Proof for Lemma 5
The proof largely follows Lemma B2, B3 in Chuang and Kim [2023]. We denote $H=\left\{H_{t}\right\}$ the natural filtration generated by the right-censored sales data, i.e $H_{t}=\sigma\left\{(Y_{i},\delta_{i}):i\leq t\right\}$, where $Y_{t}=D_{t}\wedge y_{t}$ and $\delta_{t}=1\left[D_{t}<y_{t}\right]$.
According to the proof of Lemma B2 and B3 in Chuang and Kim [2023], we have
@@
From Chuang and Kim [2023], we have
@@ A.4 Proof for Lemma 1
Since $\theta_{t}\sim\operatorname{Gamma}(\alpha_{t},\beta_{t})$, we have $\frac{1}{\theta_{t}}\sim\operatorname{InverseGamma}(\alpha_{t},\beta_{t})$. According to Chen [2014] Theorem 20, we have
@@ A.6 Proof for Lemma 11
Recall that $M_{t}=\sum_{i=0}^{t-1}\left(\delta_{i}-\mathbb{E}\left[\delta_{i}\mid\mathcal{H}_{i-1}\right]\right)$ defined in Lemma 6 and $M_{t}$ is a martingale with bounded increments (specifically, bounded by 2 ), by Azuma’s inequality we have,
@@
A.7 Proof for Lemma 8
Proof.
Setting $\epsilon=\sqrt{\frac{1}{2t}\ln\left(\frac{C}{\delta}\right)}$, then with probability at least $1-\delta$, we have
$\sup_{x}\left|(1-G(x))\left(\hat{F}_{t}(x)-F(x)\right)\right|\leq\epsilon.$
Therefore, For any $x$ with $G(x)<1$, the following holds,
$\left|\hat{F}_{t}(x)-F(x)\right|\leq\frac{\epsilon}{1-G(x)}.$
So:
$\hat{F}_{t}(x)-\frac{\epsilon}{1-G(x)}\leq F(x)\leq\hat{F}_{t}(x)+\frac{\epsilon}{1-G(x)}.$
A.8 Proof for Theorem 2
Proof.
According to definition 4 and Lemma 8 combining with assumption 1, we have
$\left|F_{\hat{\theta}_{t}}-\hat{F}_{t}\right|\leq\left|F_{\theta_{\star}}-\hat{F}_{t}\right|=\left|F_{\star}-\hat{F}_{t}\right|.$
Then we have
[…]
Given the Lipchitz assumption (2), we have for each $t\in[T]$,
$\hat{\theta}_{t}-\theta_{\star}\leq L\left|F_{\hat{\theta}_{t}}-F_{\star}\right|.$
Since the plug-in estimator $\hat{\theta}$ defined in 4 gives $\theta_{\star}$ a Minimum mean square estimation. Referring to Algorithm 1 and Using the MSE as risk, the Bayes estimate of the unknown parameter is simply the mean of the posterior distribution and is known as the minimum mean square error (MMSE) estimator.
Then, we define the newsvendor cost function parametrized by estimator $\hat{\theta}_{t}$ denoted as $g_{\hat{\theta}_{t}}(y_{t},D_{t})$. The newsvendor cost function parametrized by km estimator denoted as $\hat{g}(y_{t},D_{t})$. And the newsvendor cost function parametrized by the function class $\Theta$ denoted as $g_{\theta_{\star}}(y_{t},D_{t})$, the following inequality holds since the derivative of the newsvendor cost is Lipchitz.
$\left|g_{\hat{\theta}_{t}}(y_{t},D_{t})-\hat{g}(y_{t},D_{t})\right|\leq\left|g_{\theta_{\star}}\left(y_{\star},D_{t}\right)-\hat{g}(y_{t},D_{t})\right|=\left|g\left(y_{\star},D_{t}\right)-\hat{g}(y_{t},D_{t})\right|.$
According to Russo and Van Roy [2014], we define $U_{t}(y):=\hat{g}(y_{t},D_{t})$ and $y_{t}^{km}=\arg\min_{y}\hat{g}(y_{t},D_{t})$. Then for each $t\in[T]$, we have The per-period Bayesian regret defined in LABEL:def:bayesian_regret can be decomposed as:
$g\left(y_{t},D_{t}\right)-g(y_{\star},D_{t})\leq[g\left(y_{t},D_{t}\right)-U_{t}(y_{\star})]+[U_{t}(y_{t}^{km})-g(y_{\star},D_{t})],$
Bounding Term I: $g\left(y_{t},D_{t}\right)-U_{t}(y_{\star})$
From Lemma 8, for any $x$ such that $G(x)<1$, we have
$|\hat{F}_{t}(x)-F_{\theta_{\star}}(x)|\leq\frac{\epsilon_{t}}{1-G(x)}.$
By Lipschitz continuity of $g_{\theta}$ in distribution as stated in Assumption 2, we get:
$g\left(y_{t},D_{t}\right)-U_{t}(y_{\star})\leq|g_{\theta_{\star}}(y_{\star},D_{t})-\hat{g}(y_{\star},D_{t})|\leq\frac{L\epsilon_{t}}{1-G(y_{\star})}.$
Bounding Term II: $U_{t}(y_{t}^{km})-g(y_{\star},D_{t})$
Using Lemma 8 and the fact that cost function $U_{t}(y):=\hat{g}(y_{t},D_{t})$ is Lipschitz with respect to its first-order derivative $\hat{f}_{t}$. Then,
$U_{t}(y_{t}^{km})-g(y_{\star},D_{t})\leq\frac{(h+p)\epsilon_{t}}{1-G(y_{\star})}.$
Combining both terms, we obtain:
$g\left(y_{t},D_{t}\right)-g(y_{\star},D_{t})\leq\frac{L\epsilon_{t}}{1-G(y_{\star})}+\frac{2(h+p)\epsilon_{t}}{1-G(y_{t})}.$
Summing over $t=1$ to $T$ and taking expectations:
$\operatorname{Regret(T,\pi,\theta_{\star})}$ $=\mathbb{E}\left[\sum_{t=1}^{T}g\left(y_{t},D_{t}\right)-\sum_{t=1}^{T}g\left(y_{\star},D_{t}\right)\mid\theta_{\star}\right]$
$\leq\sum_{t=1}^{T}\mathbb{E}\left[\frac{L\epsilon_{t}}{1-G(y_{\star})}+\frac{2(h+p)\epsilon_{t}}{1-G(y_{t})}\right].$
There exists a constant $G_{\max}<1$ such that $G(x)\leq G_{\max}$ for all $x$ in the support of the demand distribution.
From Lemma 8, we have:
$\epsilon_{t}=\sqrt{\frac{1}{2t}\ln\left(\frac{1}{\delta}\right)}$
Substituting this into our regret bound:
$\operatorname{Regret}(T,\pi,\theta_{\star})$ $\leq\sum_{t=1}^{T}\mathbb{E}\left[\frac{L\epsilon_{t}}{1-G(y_{\star})}+\frac{2(h+p)\epsilon_{t}}{1-G_{\max}}\right]$
$=\sum_{t=1}^{T}\left(\frac{L}{1-G(y_{\star})}+\frac{2(h+p)}{1-G_{\max}}\right)\sqrt{\frac{1}{2t}\ln\left(\frac{1}{\delta}\right)}$
$\leq\frac{16L(h+p)}{1-G_{\max}}\sqrt{\frac{\ln(1/\delta)}{2}}\sum_{t=1}^{T}\frac{1}{\sqrt{t}}$
$\leq\frac{16L(h+p)}{1-G_{\max}}\sqrt{\ln(T)T}$
A.9 Auxiliary Lemmas
Lemma 10.
@@ A.9 Auxiliary Lemmas
$\alpha_{t}-1$ $=\alpha_{0}+\sum_{i=0}^{t-1}\left(1-e^{-\theta_{\star}y_{i}^{k}}\right)+M_{t}-1$ (8a)
$\geq\alpha_{0}-1+\left(t-1\right)\left(1-\exp\{-\theta_{\star}L^{k}\right)+M_{t}$ (8b)
$\geq t\left(1-\exp\{-\theta_{\star}L^{k}\}\right)+\alpha_{0}-1-\left(1-\exp\{-\theta_{\star}L\}^{k}\right)-\sqrt{8t}\ln{\left(\frac{2t^{2}}{\delta}\right)}$ (8c)
@@
(8a) is derived from Lemma 6. (8b) comes from the fact that $y_{t}\geq L$ for all $t$ when the event $\xi$ holds. (8c) comes from the following Lemma 11, which is proved in Appendix A.6.
Lemma 11.
@@
$\geq 1-\sum_{t=1}^{T}\frac{\delta}{T}-\sum_{t=1}^{T}\frac{\delta}{t^{2}}-\sum_{t=1}^{T}\frac{\delta}{t^{2}}-\sum_{t=1}^{T}\left(\frac{2}{e}\right)^{\alpha_{t}}$ (9a)
$\geq 1-\delta-\frac{\pi^{2}}{6}\delta-\frac{\pi^{2}}{6}\delta-\delta$ (9b)
@@
For (9a), the first term comes from Lemma 4, the second term comes from Lemma 5. The third term comes from Lemma 11. The fourth term comes from Lemma 1. (9b) comes from the following, recall $\alpha_{0}\geq\frac{\ln{\frac{T}{\delta}}}{\ln{\frac{e}{2}}}$.
@@
A.10 Additional Experiments
In this section, we conduct numerical experiments for normal distributions. To assess the impact of varying service levels, defined as $\gamma=\frac{p}{p+h}$, we set $p=1$ and adjust $h$ to achieve service levels of 50 %, 90%, and 98 %. All policies are evaluated with prior parameters $\alpha_{0}=\beta_{0}=4$ and time horizon $T=600$. Each algorithm runs across 100 independent trials, with average cumulative regret reported. The results are presented in two figures: (1) a performance comparison of TS, UCB, and OCO algorithms (Figure 3), and (2) a comparison of TS and the myopic policy relative to optimal cost (Figure 4). Across all service levels, TS demonstrates superior performance compared to benchmark methods and achieves faster convergence than the myopic policy, demonstrating its effectiveness in balancing exploration and exploitation.
Figure 3: (Normal Distribution) Compare TS with OCO and UCB
Figure 4: (Normal Distribution) Compare TS with Myopic